# Treatment heterogeneity with right-censored outcomes using grf

### 05 Dec 2023 v1

@@
We use data from the National Job Training Partnership Act (JPTA) study that assessed the impact of eligibility for an employment training program on employment outcomes using a randomized controlled trial. All participants were interviewed at enrollment before randomization; they were then re-interviewed sometime between one and three years post-enrollment. The study enrolled participants who were unemployed at the time, and one natural question to ask using data from this study is: How did being randomized into treatment affect the amount of time it took for study participants to find a job (Crommen et al., 2022, Frandsen, 2015)? This time-to-event outcome is subject to random censoring because not all participants had yet found a new job at the time of the second survey (which was administered with different delays post-enrollment for different participants); however, the censoring appears plausibly ignorable here and so causal survival forests can be used to correct for it.
@@
we see that most point estimates for $\hat{\tau}(X_{i})$ are large and positive, suggesting the program offering reduces the unemployment duration, and that the magnitude of the effect varies across individuals.
@@
Given an estimated function $\hat{\tau}(\cdot)$ that appears to detect meaningful heterogeneity, there are a variety of ways we can probe further into how this heterogeneity manifests itself via different covariate characteristics $X_{i}$, by tabulating various summary statistics, such as perhaps looking at the distributions of different covariates over low and high predicted CATEs. Here we simply conclude with a brief example of showing the covariate means for the full sample, and the covariate mean for the 20% of units with the largest predicted decrease in unemployment duration, which appears to indicate there are a larger fraction of parents and married people in the group that is predicted to have a large effect.

### 21 Jan 2024 v2

@@
We use data from the National Job Training Partnership Act (JPTA) study that assessed the impact of eligibility for an employment training program on employment outcomes using a randomized controlled trial. All participants were interviewed at enrollment before randomization; they were then re-interviewed sometime between one and three years post-enrollment. The study enrolled participants who were unemployed at the time, and one natural question to ask using data from this study is: How did being randomized into treatment affect the amount of time it took for study participants to find a job (Crommen et al., 2023, Frandsen, 2015)? This time-to-event outcome is subject to random censoring because not all participants had yet found a new job at the time of the second survey (which was administered with different delays post-enrollment for different participants); however, the censoring appears plausibly ignorable here and so causal survival forests can be used to correct for it.
@@
If we look at the distribution of the CATE estimates we see that most point estimates for $\hat{\tau}(X_{i})$ are large and positive, suggesting the program offering reduces the unemployment duration, and that the magnitude of the effect varies across individuals.
@@
Given an estimated function $\hat{\tau}(\cdot)$ that appears to detect meaningful heterogeneity, there are a variety of ways we can probe further into how this heterogeneity can be described using available covariates $X_{i}$. One simple option is to examine the distributions of covariates across subregions defined in terms of quantiles of the estimated CATEs. Here we simply conclude with a brief example of showing the covariate means for the full sample, and the covariate mean for the 20% of units with the largest predicted decrease in unemployment duration, which appears to indicate there are a larger fraction of parents and married people in the group that is predicted to have a large effect.